graph linear equation using the slope intertcept method of the following
5(x-3)=3y+6

Answer by mananth(16949)

5(x-3)=3y+6
5x-15=3y+6
5x-15-6 = 3y
5x-21=3y
y=(5/3)x -7
find x & y intercept
y intercept plug x=0
y=-7 ....... (0,-7)
x intercept plug y =0
(5/3)x=7
x=21/5............ (21/5, 0)